OpenStudy (anonymous):

Please assist me? I'll answer anything related to English, Psychology, or Philosophy in return? ;) Solve 2x + 5y = 10 for y. a.) y= 2x + 2 b.) y= -2x + 10 c.) y= - 2/5x + 2 d.) y= 2/5 x + 10

OpenStudy (anonymous):

i think its b)

OpenStudy (anonymous):

Not "C"?

OpenStudy (anonymous):

It's either b or C

OpenStudy (anonymous):

sorry nooo

OpenStudy (anonymous):

You need to get y on its own, so start by getting rid of the 2x from the left hand side. i.e, subtract 2x from both sides. You now have: \[5y=10-2x\] Now, you want to find out what y alone is, not what 5y is. So divide both sides by 5 to get: \[y=2-\frac{2x}{5}\]

OpenStudy (anonymous):

2x + 5y = 10 -> 5y = 10 - 2x => y = ....?

OpenStudy (anonymous):

It's "C"

OpenStudy (anonymous):

Greekstud, where's your proof? It's C, I believe

OpenStudy (anonymous):

Does that make sense to you @InsanelyChaotic ? When solving an equation for a particular variable, you need to perform operations to both sides of the equation to get the variable (y) on its own.

OpenStudy (anonymous):

wait.... i really did a mistake its a) because 5y=10-2x y=10/5 -2x y=2-2x

OpenStudy (anonymous):

@Traxter Yes, that does make sense

OpenStudy (anonymous):

Excellent :)

OpenStudy (anonymous):

WAIT, why would you subtract and not divide?

OpenStudy (anonymous):

Why would you subtract what?

OpenStudy (anonymous):

Nevermind, you divided by 5 on both sides

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Thank you kind sir

OpenStudy (anonymous):

You're very welcome.

OpenStudy (anonymous):

I'm officially a fan of yours